OpenStudy (dveliz26):

What net force is required to bring a 1500 kg car to rest from a speed of 100 km/h over a distance of 55 m?

OpenStudy (osprey):

This is one of the KINEMATIC EQUATIONS applied, for once, to DECLERATION, and INCLUDING a force. So, it ALSO "wants" F=ma, Newton's second law of motion, albeit in formula form (N2 goes on about rates of change of MOMENTUM) v squared = u squared minus 2 Force/mass s. v squared = v = 0. 0 =[100x10^3/(60x60)]^2-Force/1500 x 55 If my scribbling is right, solve for Force to get ... force. Bon voyage et bon chance http://perendis.webs.com

OpenStudy (hoot_hoot):

F = ma Mass is given. Acceleration is not, but you can solve for it using one of the kinematic equations, you can use the: \[v_{f}^{2} = v_{i}^{2} + 2ad\] (final velocity)^2 = (initial velocity)^2 + 2(acceleration)(distance) final velocity is 100 km/h. Initial velocity is 0 km/h because of the word "rest." Distance is 55 m. You can derive the acceleration from the formula \[v_{f}^{2} = 2ad\] \[a = \frac{ v_{f}^{2} }{ 2d } = \frac{ 10 000 \frac{ km }{ h } }{ 110 } = ?\] Multiply the acceleration by the mass (1 500 kg) to get the force.

OpenStudy (hoot_hoot):

It must be: \[a = \frac{ 10 000 \frac{ km^ {2} }{ h^ {2}} }{ 110 m }\] not \[a = \frac{ 10 000 \frac{ km }{ h} }{ 110 }\] You can convert first 110 m to km. \[110 m \times \frac{ 1 km }{ 1 000 m } = ?\] and then divide it to 10 000 km^2/h^2.

OpenStudy (hoot_hoot):

From that, you can now find the net force.